A company that manufactures purgable hydrogen sulfide monitors will make deposits such that each one is 5% larger than the preceding one. How large must the first deposit at the end of year 1 be if the deposits extend through year 10 and the 5th deposit is $4750? Use an interest rate of 6% per year The first deposit at the end of year 1 must be$
Answer #1

Deposit on year 5 = 4750
Growth = 6%
Deposit on year 1 = 4750/(1+5%)5 = 4750/1.055 = 3721.75

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well